Glucose has the formula C6H12O6. What is the maximum mass of carbon that could be obtained from 200.00g of glucose?
1
Expert's answer
2020-05-03T15:12:27-0400

Solution:

Glucose has the formula C6H12O6.

This means every mole of glucose contains 6 moles of carbon.

To convert our mass of glucose to moles of glucose:

Moles of glucose = Mass of glucose / Molar mass of glucose

The molar mass of C6H12O6 is 180.156 g/mol.

Moles of glucose = (200.00 g) / (180.156 g/mol) = 1.11 mol


One mole of glucose contains 6 moles of C, therefore:

Moles of carbon = (1.11 mol C6H12O6) × (6 mol of C / 1 mol C6H12O6) = 6.66 mol of carbon (C).


Now, using the molar mass of carbon (12.0107 g/mole) to calculate the total mass of carbon in the sample.

Mass of carbon = (12.0107 g/mol) × (6.66 mol) = 79.99 g = 80.00 g.

The maximum mass of carbon is 80.00 g.


Answer: 80.00 g is the maximum mass of carbon that could be obtained from 200.00g of glucose.